OpenStudy (ronaldo7):

I need to do this question I just need to know is this formula is it d=r/t correct and the "d" represnts distance and the "r" represents rate, and the "t" represents time, and the question is two towns are 460miles apart. A car leaves the first town traveling toward the second town at 60mph. At the same time, a motorcuclist leaves the other town and heads toward the first town at 55mph. How long will it take for the drivers to meet.

OpenStudy (abb0t):

Rerrange your formula and solve for "t". It's asking, "how long" meaning time, I believe in hours. As for distance, use ∆d, where ∆ is the change (or difference)

OpenStudy (ronaldo7):

So what is the formula.

OpenStudy (whpalmer4):

d = r * t You can determine this by unit analysis. rate is length/time, distance is length, time is time. d = r * t -> length/time * time = length

OpenStudy (whpalmer4):

You that the towns are 460 miles apart. When the drivers meet, they will (between them) have covered all 460 miles. One driver goes 60 mph, the other 55 mph. Together, they use up 60+55 = 115 mph. How many hours does it take to cover 460 miles at 115 mph?

OpenStudy (whpalmer4):

A slightly more complicated problem (1 more step) would be to ask where the two drivers met. Another twist might be to have one driver start before the other, in which case you would subtract off the distance covered in the "head start" period from the total distance the two have to cover, then proceed as before.

OpenStudy (anonymous):

\[R=speed=velocity=\frac{\Delta D}{\Delta T}\] however you started with the wrong equation Speed or velocity = distance over time \[S=\frac{D}{T}=R\]